40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Aug 25 11:02:39 2009, rana, Summary, PSL, Temperature Box D980400-B.pdf
    Reply  Tue Aug 25 18:42:42 2009, steve, Update, PSL, reference cavity temp box temporarly out of order 
       Reply  Tue Aug 25 21:36:28 2009, Alberto, Update, PSL, reference cavity temp box temporarly out of order 2009-08-25_PSLtrend.png
       Reply  Tue Aug 25 21:55:11 2009, rana, Update, PSL, reference cavity temp box temporarly out of order Untitled.pngrc-heater.jpg
          Reply  Wed Aug 26 15:42:17 2009, Alberto, Update, PSL, reference cavity temp box temporarly out of order 
             Reply  Wed Aug 26 16:11:41 2009, Alberto, Update, PSL, reference cavity temp box temporarly out of order 
                Reply  Wed Aug 26 16:31:34 2009, steve, Update, PSL, reference cavity temp box temporarly out of order 
Message ID: 1949     Entry time: Wed Aug 26 15:42:17 2009     In reply to: 1946     Reply to this: 1951
Author: Alberto 
Type: Update 
Category: PSL 
Subject: reference cavity temp box temporarly out of order 

Quote:

There's no elog entry about what work has gone on today, but it looks like Peter took apart the reference cavity temperature control around 2PM.

I touched the reference cavity by putting my finger up underneath its sweater and it was nearly too hot to keep my finger in there. I looked at the heater power supply front panel and it seems that it was railed at 30 V and 3 A. The nominal value according to the sticker on the front is 11.5 V and 1 A.

So I turned down the current on the front panel and then switched it off. Otherwise, it would take it a couple of days to cool down once we get the temperature box back in. So for tonight there will definitely be no locking. The original settings are in the attached photo. We should turn this back on with its 1A setting in the morning before Peter starts so that the RC is at a stable temp by the evening. Its important NOT to turn it back on and let it just rail. Use the current limit to set it to 1 A. After the temperature box is back in the current limit can be turned back up to 2A or so. We never need the range for 3A, don't know why anyone set it so high.

While Peter King is still working on the reference cavity temperature box, I turned the power supply for the reference cavity's heater back on. Rana turned it off last night since the ref cav temperature box had been removed.

I just switched it on and turned the current knob in the front panel until current and voltage got back to their values as in Rana's picture.

I plan to leave it like that for half an hour so that the the cavity starts warming up. After that, I'll turn the current back to the nominal value as indicated in the front panel.

ELOG V3.1.3-